# Missing Data Imputation by Reducing Mutual Information with Rectified Flows

### v2

@@ Appendix D Proof of Proposition 2
where $\mathrm{const.}$ does not depend on $\mathbf{g}$. The KL divergence $\mathrm{D}[p_{\mathbf{g}}(\mathbf{x}_{1-\mathbf{m}}|\mathbf{x}_{\mathbf{m}},\mathbf{m})|q(\mathbf{x}_{1-\mathbf{m}}|\mathbf{x}_{\mathbf{m}})]$ is zero if and only if $p_{\mathbf{g}}(\mathbf{x}_{1-\mathbf{m}}|\mathbf{x}_{\mathbf{m}},\mathbf{m})=q(\mathbf{x}_{1-\mathbf{m}}|\mathbf{x}_{\mathbf{m}})$. Assuming $p(\mathbf{x}_{\mathbf{m}}|\mathbf{m})$ is positive everywhere, $A_{\mathbf{g}}$ is minimized if and only if $\mathrm{D}[p_{\mathbf{g}}(\mathbf{x}_{1-\mathbf{m}}|\mathbf{x}_{\mathbf{m}},\mathbf{m})|q(\mathbf{x}_{1-\mathbf{m}}|\mathbf{x}_{\mathbf{m}})]$ is zero. Thus, if $p(\mathbf{m})$ is strictly positive, $\mathrm{D}\left[\mathbb{P}_{\mathbf{X}^{(t)}(\mathbf{g}),\mathbf{M}}|\mathbb{P}_{\mathbf{X}^{(t-1)}}\otimes\mathbb{P}_{\mathbf{M}}\right]$ is minimized if and only if $p_{\mathbf{g}}(\mathbf{x}_{1-\mathbf{m}}|\mathbf{x}_{\mathbf{m}},\mathbf{m})=q(\mathbf{x}_{1-\mathbf{m}}|\mathbf{x}_{\mathbf{m}})$.
Appendix E Experimental Setup
@@ E.1 Datasets
[…]
@@
We randomly sample 5 000 32$\times$32 images and apply pixel-level MCAR masks at varying rates. All three colour channels of each missing pixel are masked.
@@
We randomly sample 5 000 64$\times$64 images and apply channel-level MCAR masks at varying rates. All three colour channels of each missing pixel are masked independently.
E.2 Evaluation Protocol
@@ E.4 Computational Resources
E.5 Baseline Implementations
We evaluate HyperImpute and MICE using the official HyperImpute repository with default settings. [^2] https://github.com/vanderschaarlab/hyperimpute.
We adapt the official implementations of KnewImp [^3] https://github.com/JustusvLiebig/NewImp., TabCSDI [^4] https://github.com/pfnet-research/TabCSDI. and GAIN [^5] https://github.com/jsyoon0823/GAIN. from their respective GitHub repositories. MissDiff is reimplemented based on the algorithm described in the original publication [Ouyang et al., 2023]. All baselines use default hyperparameters unless otherwise stated.
Appendix F Additional Experimental Results
F.1 Performance on Synthetic Data
@@
Overall, HyperImpute is optimal when minimizing per-entry error, achieving roughly 8% lower RMSE and 14% lower MAE than MIRI. In contrast, MIRI reduces MMD by more than 60% relative to HyperImpute, thereby better preserving the underlying data distribution despite a modest increase in point-wise error. MissDiff offers a balanced trade-off, whereas GAIN, TabCSDI, and KnewImp underperform on both fronts.
F.2 Additional UCI Regression Experiments
Figure 4 shows the MMD results on ten UCI regression benchmarks under three MCAR levels: 20% (top), 40% (middle), and 60% (bottom). Lower values indicate better distributional matching.
20% Missingness.
MIRI attains an average MMD of $\approx 0.05$, compared to HyperImpute ($\approx 0.07$), MissDiff ($\approx 0.04$), and MICE ($\approx 0.10$). KnewImp, TabCSDI and GAIN suffer much higher MMD—$\approx 0.38$, $\approx 0.54$, and $\approx 0.60$, respectively. The largest relative gains of MIRI over HyperImpute occur on high-dimensional datasets (gas, pumadyn32nm), where MIRI reduces MMD by over $\approx 30\%$.
40% Missingness.
MIRI’s MMD increases slightly to $\approx 0.04$, while HyperImpute remains at $\approx 0.06$ and MissDiff at $\approx 0.03$. MICE degrades to $\approx 0.07$. KnewImp, TabCSDI and GAIN remain elevated at $\approx 0.27$, $\approx 0.55$, and $\approx 0.55$, respectively.
60% Missingness.
Under severe missingness, MIRI’s MMD is $\approx 0.05$, still below HyperImpute ($\approx 0.07$) and MissDiff ($\approx 0.04$). MICE further deteriorates to $\approx 0.10$, while KnewImp, TabCSDI and GAIN exhibit persistently high MMD ($\approx 0.21$, $\approx 0.55$, and $\approx 0.52$, respectively).
@@ F.2 Additional UCI Regression Experiments
These results confirm that MIRI consistently delivers strong distributional fidelity, remains robust under high missing rates, and scales effectively to high-dimensional datasets.
@@
Figure 4: MMD on 10 UCI datasets (Above: 20% missingness, Middle: 40% missingness, Below: 60 % missingness). The lower the better.
F.3 Additional CIFAR-10 Experiments
@@
We evaluate the imputation quality of MIRI, GAIN, KnewImp, and HyperImpute on 15 randomly selected CIFAR-10 images corrupted under pixel-level MCAR with missing rates of 20%, 40% and 60% (see Figures 5, 6 and 7). Across all missing-data regimes, MIRI consistently produces accurate, visually coherent reconstructions, in agreement with the quantitative metrics in Table 1. Although HyperImpute achieves competitive performance at 20% missingness, MIRI outperforms it at both 40% and 60% missingness.
@@
Figure 5: 15 uncurated 32$\times$32 CIFAR-10 images and their imputations. 20% of pixels randomly removed from all RGB channels.
@@
Figure 6: 15 uncurated 32$\times$32 CIFAR-10 images and their imputations. 40% of pixels randomly removed from all RGB channels.
@@
Figure 7: 15 uncurated 32$\times$32 CIFAR-10 images and their imputations. 60% of pixels randomly removed from all RGB channels.
Appendix G Illustration of Sequential Imputation
Figure 8 illustrates the sequential imputation process of MIRI.
Figure 8: A basic schematic of the sequential imputation algorithm for $t=1\dots T$. $\left[\mathbf{X}\right]_{i}$ denotes the $i$-th components of $\mathbf{X}$. In this example, $\mathbf{X}^{(0)}$ are initialized using the mean ($\bar{\mathbf{X}}$) imputation.
Appendix H Euler ODE Solver
Algorithm 2 ODE Solver with Euler Method
@@ Appendix H Euler ODE Solver
0:  Velocity field $\mathbf{v}$, Initial condition $\{(\mathbf{X}_{i},\mathbf{M}_{i})\}$, Number of Euler steps $N$.
1:  for $k=1$ to $N$ do

### v3

@@ Appendix E Proof of Proposition 2
Thus, if $p(\mathbf{m})$ is strictly positive, $\mathrm{D}\left[\mathbb{P}_{\mathbf{X}^{(t)}(\mathbf{g}),\mathbf{M}}\|\mathbb{P}_{\mathbf{X}^{(t-1)}}\otimes\mathbb{P}_{\mathbf{M}}\right]$ is minimized if and only if $p_{\mathbf{g}}(\mathbf{x}_{1-\mathbf{m}}|\mathbf{x}_{\mathbf{m}},\mathbf{m})=q(\mathbf{x}_{1-\mathbf{m}}|\mathbf{x}_{\mathbf{m}})$.
Appendix F Validity of MIRI under MAR Setting
Recall that we denote the observed components of a vector $\mathbf{X}$ as $\mathbf{X}_{\mathbf{M}}$ and the missing components as $\mathbf{X}_{1-\mathbf{M}}$. Under the Missing at Random (MAR) assumption, the goal is to minimize the conditional mutual information:
$\mathrm{I}[\mathbf{X}_{1-\mathbf{M}}(\mathbf{g});\mathbf{M}|\mathbf{X}_{\mathbf{M}}]=\mathbb{E}_{(\mathbf{x},\mathbf{m})\sim(\mathbf{X},\mathbf{M})}\left[\log\frac{p_{\mathbf{g}}(\mathbf{x}_{1-\mathbf{m}},\mathbf{m}|\mathbf{x}_{\mathbf{m}})}{p_{\mathbf{g}}(\mathbf{x}_{1-\mathbf{m}}|\mathbf{x}_{\mathbf{m}})p(\mathbf{m}|\mathbf{x}_{\mathbf{m}})}\right].$ (56)
Following a similar argument to the proof of Proposition 1, we can define an iterative algorithm to reduce this conditional mutual information. At each iteration $t$, we find the optimal imputer $\mathbf{g}_{t}$ by solving:
$\mathbf{g}_{t}\in\arg\min_{\mathbf{g}}\mathbb{E}_{(\mathbf{x},\mathbf{m})\sim(\mathbf{X}_{\mathbf{g}}^{(t)},\mathbf{M})}\left[\log\frac{p_{\mathbf{g}}(\mathbf{x}_{1-\mathbf{m}},\mathbf{m}|\mathbf{x}_{\mathbf{m}})}{q(\mathbf{x}_{1-\mathbf{m}}|\mathbf{x}_{\mathbf{m}})q(\mathbf{m}|\mathbf{x}_{\mathbf{m}})}\right],$ (57)
where $p_{\mathbf{g}}$ is the density corresponding to the distribution of $(\mathbf{X}_{\mathbf{g}}^{(t)},\mathbf{M})$, and $q$ is the density of $\mathbf{X}^{(t-1)}$ from the previous iteration.
We now show that $p_{\mathbf{g}}(\mathbf{x}_{1-\mathbf{m}}|\mathbf{m},\mathbf{x}_{\mathbf{m}})=q(\mathbf{x}_{1-\mathbf{m}}|\mathbf{x}_{\mathbf{m}})$ is a sufficient condition for $\mathbf{g}$ to be optimal. Analogous to the proof of Proposition 2, we can rewrite the objective function. Since the missingness mechanism is fixed and does not depend on the imputer $\mathbf{g}$, we have $p(\mathbf{m}|\mathbf{x}_{\mathbf{m}})=q(\mathbf{m}|\mathbf{x}_{\mathbf{m}})$. The objective then simplifies:
$\mathbb{E}_{(\mathbf{x},\mathbf{m})\sim(\mathbf{X}_{\mathbf{g}}^{(t)},\mathbf{M})}\left[\log\frac{p_{\mathbf{g}}(\mathbf{x}_{1-\mathbf{m}},\mathbf{m}|\mathbf{x}_{\mathbf{m}})}{q(\mathbf{x}_{1-\mathbf{m}}|\mathbf{x}_{\mathbf{m}})q(\mathbf{m}|\mathbf{x}_{\mathbf{m}})}\right]$ (58)
$=$ $\mathbb{E}_{(\mathbf{x},\mathbf{m})\sim(\mathbf{X}_{\mathbf{g}}^{(t)},\mathbf{M})}\left[\log\frac{p_{\mathbf{g}}(\mathbf{x}_{1-\mathbf{m}}|\mathbf{m},\mathbf{x}_{\mathbf{m}})p(\mathbf{m}|\mathbf{x}_{\mathbf{m}})}{q(\mathbf{x}_{1-\mathbf{m}}|\mathbf{x}_{\mathbf{m}})q(\mathbf{m}|\mathbf{x}_{\mathbf{m}})}\right]$ (59)
$=$ $\mathbb{E}_{(\mathbf{x},\mathbf{m})\sim(\mathbf{X}_{\mathbf{g}}^{(t)},\mathbf{M})}\left[\log\frac{p_{\mathbf{g}}(\mathbf{x}_{1-\mathbf{m}}|\mathbf{m},\mathbf{x}_{\mathbf{m}})}{q(\mathbf{x}_{1-\mathbf{m}}|\mathbf{x}_{\mathbf{m}})}\right]$ (60)
$=$ $\mathbb{E}_{(\mathbf{m},\mathbf{x}_{\mathbf{m}})\sim(\mathbf{M},\mathbf{X}_{\mathbf{M}})}\mathbb{E}_{\mathbf{x}_{1-\mathbf{m}}\sim p_{\mathbf{g}}(\cdot|\mathbf{m},\mathbf{x}_{\mathbf{m}})}\left[\log\frac{p_{\mathbf{g}}(\mathbf{x}_{1-\mathbf{m}}|\mathbf{m},\mathbf{x}_{\mathbf{m}})}{q(\mathbf{x}_{1-\mathbf{m}}|\mathbf{x}_{\mathbf{m}})}\right]$ (61)
$=$ $\mathbb{E}_{(\mathbf{m},\mathbf{x}_{\mathbf{m}})\sim(\mathbf{M},\mathbf{X}_{\mathbf{M}})}\mathrm{D}[p_{\mathbf{g}}(\cdot|\mathbf{m},\mathbf{x}_{\mathbf{m}})\|q(\cdot|\mathbf{x}_{\mathbf{m}})].$ (62)
The inner expectation is the KL divergence between the conditional distributions $p_{\mathbf{g}}(\mathbf{x}_{1-\mathbf{m}}|\mathbf{m},\mathbf{x}_{\mathbf{m}})$ and $q(\mathbf{x}_{1-\mathbf{m}}|\mathbf{x}_{\mathbf{m}})$. The KL divergence is non-negative and is minimized (to zero) if and only if the two distributions are equal:
$p_{\mathbf{g}}(\mathbf{x}_{1-\mathbf{m}}|\mathbf{m},\mathbf{x}_{\mathbf{m}})=q(\mathbf{x}_{1-\mathbf{m}}|\mathbf{x}_{\mathbf{m}}).$ (63)
This demonstrates that the same optimality condition derived for the MCAR setting in Proposition 2 also holds under the MAR assumption.
Appendix G Experimental Setup
@@ G.1 Datasets
We generate N=6  [^0] N=6\,000 samples in $\mathbb{R}^{2}$ by drawing two equally‐sized clusters of n=3 000n=3\,000 points each from isotropic Gaussians. One cluster is centered at (2,2)(-2,-2) and the other at $(2,2)$, both with standard deviation $\sigma=0.5$.
@@
We randomly sample 5 000 $\times32$ RGB images and apply pixel-level MCAR masks at varying rates. All three colour channels of each missing pixel are masked.
@@
We randomly sample 5 000 $\times64$ RGB images and apply channel-level MCAR masks at varying rates. All three colour channels of each missing pixel are masked independently.
G.2 Evaluation Protocol
@@ G.4 Computational Resources
G.5 Baseline Implementations
We evaluate HyperImpute, MICE, MIWAE, and Sinkhorn using the official HyperImpute repository with default settings. [^1] https://github.com/vanderschaarlab/hyperimpute.
We adapt the official implementations of KnewImp [^2] https://github.com/JustusvLiebig/NewImp., TabCSDI [^3] https://github.com/pfnet-research/TabCSDI., GAIN [^4] https://github.com/jsyoon0823/GAIN., and TDM [^5] https://github.com/hezgit/TDM. from their respective GitHub repositories. MissDiff is reimplemented based on the algorithm described in the original publication (ouyang_missdiff_2023). All baselines use default hyperparameters unless otherwise stated.
Appendix H Additional Experimental Results
H.1 Performance on Synthetic Data
@@
Overall, HyperImpute is optimal when minimizing per‐entry error, achieving roughly 8% lower RMSE and 14% lower MAE than MIRI. In contrast, MIRI reduces MMD by more than 60% relative to HyperImpute, thereby better preserving the underlying data distribution despite a modest increase in point‐wise error. MissDiff offers a balanced trade‐off, whereas GAIN, TabCSDI, and KnewImp underperform on both fronts.
H.2 Additional UCI Regression Experiments
@@
We comprehensively evaluate imputation performance on ten UCI regression benchmarks under MCAR, MAR, and MNAR settings (Figures 5, 6, and 7). In some high-dimensional or high-missingness scenarios, certain baselines failed; for example, HyperImpute produced runtime errors and MIWAE encountered out-of-memory (OOM) errors. The results confirm that MIRI consistently delivers strong distributional fidelity, remains robust under high missing rates, and scales effectively to high-dimensional datasets.
@@
Figure 5: MCAR MMD on 10 UCI datasets (Above: 20% missingness, Middle: 40% missingness, Below: 60 % missingness). The lower the better.
Figure 6: MAR MMD on 10 UCI datasets (Above: 40% missingness, Below: 80 % missingness). The lower the better.
Figure 7: MNAR MMD on 10 UCI datasets (Above: 20% missingness, Middle: 40% missingness, Below: 60 % missingness). The lower the better.
H.3 Additional CIFAR-10 Experiments
@@
We evaluate the imputation quality of MIRI, GAIN, KnewImp, and HyperImpute on 15 randomly selected CIFAR-10 images corrupted under pixel-level MCAR with missing rates of 20%, 40% and 60% (see Figures 8(a), 8(b) and 8(c)). Across all missing-data regimes, MIRI consistently produces accurate, visually coherent reconstructions, in agreement with the quantitative metrics in Table 2. Although HyperImpute achieves competitive performance at 20% missingness, MIRI outperforms it at both 40% and 60% missingness.
@@
(a) 15 uncurated $\times32$ CIFAR-10 images and their imputations. 20% of pixels randomly removed from all RGB channels.
@@
(b) 15 uncurated $\times32$ CIFAR-10 images and their imputations. 40% of pixels randomly removed from all RGB channels.
@@
(c) 15 uncurated $\times32$ CIFAR-10 images and their imputations. 60% of pixels randomly removed from all RGB channels.
Appendix I Computational Time
Computational cost matters in high dimensions. MIRI is end-to-end vectorized (“vector in, vector out”) and exploits parallel hardware common in image/tabular pipelines, whereas round-robin methods such as MICE and HyperImpute (Jarrett2022HyperImpute; vanbuuren2011mice) are inherently sequential and scale poorly with dimensionality.
Wall-clock runtimes.
Table 6 reports mean $\pm$ std hours to impute $1000$ samples across dimensionalities. We include diffusion-style baselines (DiffPuter (zhang2025diffputer), TabCSDI (zheng_diffusion_2023)), round-robin (HyperImpute), distribution-matching (MOT (muzellec2020missing), TDM (zhao2023transformed)) and VAE (MIWAE (mattei2019miwae)). MIRI is on par with diffusion-based methods and substantially faster than round-robin in high dimensions. On CIFAR-10 (5k images, 60% missing), HyperImpute did not finish within a 24h budget on our PyTorch+CUDA setup, while MIRI finished in $\approx$3.5h on the same hardware.
[…]
Method 50d 200d 500d 1000d 2000d 5000d
HyperImpute $0.08\pm 0.03$ $0.23\pm 0.02$ $1.18\pm 0.51$ $3.84\pm 0.99$ $11.65\pm 3.34$ $57.54\pm 0.00$
TabCSDI $0.18\pm 0.00$ $0.20\pm 0.01$ $0.26\pm 0.00$ $0.36\pm 0.00$ $0.53\pm 0.01$ $1.10\pm 0.01$
DiffPuter $2.30\pm 0.01$ $3.22\pm 0.18$ $5.06\pm 0.01$ $7.70\pm 0.05$ $12.40\pm 0.21$ $29.10\pm 0.00$
MIWAE $0.11\pm 0.00$ $0.30\pm 0.00$ $0.90\pm 0.03$ $2.02\pm 0.01$ OOM OOM
MOT $1\pm 0.00 0.33±0$$3\pm 0.00$ $0.36\pm 0.00$ $0.38\pm 0.00$ $0.49\pm 0.02$ $0.56\pm 0.00$
TDM $0.32\pm 0.00$ $0.55\pm 0.00$ $1.30\pm 0.02$ $3.30\pm 0.04$ $13.90\pm 0.83$ OOM
MIRI (Ours*) $0.17\pm 0.00$ $0.21\pm 0.00$ $0.28\pm 0.02$ $0.37\pm 0.02$ $0.59\pm 0.02$ $1.30\pm 0.01$
Appendix J Euler ODE Solver
@@ Appendix I Computational Time
0: Velocity field $\mathbf{v}$, Initial condition $\{(\mathbf{X}_{i},\mathbf{M}_{i})\}$, Number of Euler steps NN.
1:for $k=1$ to NN do